OpenStudy (anonymous):

A polynomial function can be written as (x + 2)(x + 3)(x - 5). What are the x-intercepts of the graph of this function?

OpenStudy (imstuck):

Thee x intercepts occur where x = 0. Your first factor of x+2 = 0 says that x = -2. That is an x intercept. The next one of x + 3 = 0 tells us that x = -3. That's another x intercept. The last one, x - 5 = 0 tells us that x = 5. That's the third x intercept.

OpenStudy (imstuck):

x intercepts are -2, -3, 5

OpenStudy (anonymous):

ty

OpenStudy (imstuck):

YW!!!

OpenStudy (anonymous):

one more?

OpenStudy (imstuck):

sure!

OpenStudy (imstuck):

BRB, ok?